# TomoGraphView: 3D Medical Image Classification with Omnidirectional Slice Representations and Graph Neural Networks

### v1

@@ 4.3 Volume Slicing Strategy Benchmarking
Incorporating additional complementary perspectives, as in the 2.5D slicing strategy, leads to a clear improvement in performance. For example, the average AUROC increases from 0.7335 for the 2D-axial+ configuration (three slices) to 0.7681 for the 2.5D setup (three slices) across six different oncology datasets. This trend extends to the 2.5D+ variant, which consistently outperforms its 2D-axial+ counterpart for the same number of views. While the performance differences are more pronounced between $3$, 88, and $16$ views, the gap narrows at $24$ views, suggesting that incorporating diverse perspectives is particularly beneficial when only a limited number of slices are available.
@@
Beyond canonical plane slicing, our proposed omnidirectional slicing strategy, which integrates both canonical and non-canonical orientations, further improves downstream performance. On average, across all evaluated configurations, the omnidirectional approach consistently outperforms both the 2D-axial+ and 2.5D+ methods. Furthermore, for our volumetric slicing strategy, we observe a general trend across datasets. Increasing the number of views typically leads to higher performance. This relationship remains consistent for the omnidirectional approach, with the sole exception of the kidney dataset, highlighting a clear positive correlation between the number of views and overall model performance. Furthermore, although the proposed omnidirectional slicing strategy does not achieve the highest performance on every individual dataset, it consistently surpasses its 2D-axial+ and 2.5D+ counterparts for the same number of views when considering the mean performance across all datasets. This demonstrates the method’s robust and, in several cases, superior capability in capturing complementary volumetric information relevant for downstream classification.
4.4 Performance Evaluation of TomoGraphView
@@
Building on the benchmarking results presented in Section 4.3, our experiments demonstrate that the proposed omnidirectional slicing approach provides a robust and effective alternative to conventional volume slicing strategies, surpassing their capabilities across datasets and, remarkably, resulting in a higher overall average downstream performance. However, since the MLP prediction head is inadequate to account for relationships between omnidirectional slice encodings, we shift our focus toward aggregating slice-level representations using the proposed spherical mesh-based graph topology (see Section 3.3) and its corresponding graph-based feature aggregation mechanism. In this experiment, we compare Omnidirectional Volume Slicing + MLP (Section 4.3) against our proposed TomoGraphView framework (i.e., Omnidirectional Volume Slicing + GNN). Both methods operate on identical omnidirectional slice inputs, ensuring that any observed performance differences arise solely from the aggregation strategy rather than the input representations. The corresponding results are summarized in Table 2.
The key distinction lies in the GNN’s ability to explicitly integrate both local and global information by propagating slice-level features across nodes through edge connections defined by the graph structure (see Section 3.3). This mechanism is not available to the MLP, which instead processes all slice information through fully connected layers without explicit relational modeling. Furthermore, given that the graph topology approximates a three-dimensional geometric structure (i.e., a sphere), we argue that a graph-based representation is better suited to preserve the inherent spatial relationships among slices compared to dense connections in an MLP. Consequently, the graph structure provides a more faithful representation of the 33D volume when represented through 22D slices and their corresponding relative position.
@@
The advantage of propagating information across slices via a meaningful underlying graph structure is supported by the results in Table 2. Specifically, when comparing the average performance of the Omnidirectional Volume Slicing + MLP against our TomoGraphView framework (Omnidirectional Volume Slicing + GNN) across all six datasets, our proposed framework consistently outperforms the MLP for every view configuration, ultimately increasing from $0.8154$ to $0.8372$ in AUROC performance for the best performing model in each case. Notably, our TomoGraphView framework achieves superior performance across nearly all dataset–view combinations, with the exception of Head-Neck at $8$ views and Soft-Tissue at $24$ views. These findings strongly highlight the effectiveness of the graph-based approach over the MLP baseline.
Table 2: Quantitative Area Under the Receiver Operating Characteristic Curve (AUROC) results for the performance comparison between omnidirectional volume slicing in combination with an MLP against our proposed TomoGraphView framework. Evaluation is performed across different datasets and view configurations. The rightmost column shows the average performance from each method and view. The best performing method on a per dataset basis is indicated in bold.
Method Views Breast Brain Head-Neck Kidney Liver Soft-Tissue Mean
Omnidirectional Volume Slicing + MLP 8 0.6627 0.9457 0.6790 0.7306 0.9372 0.7702 0.7876
16 0.6689 0.9491 0.6999 0.7450 0.9352 0.8090 0.8012
24 0.6794 0.9492 0.7384 0.7229 0.9421 0.8602 0.8154
Omnidirectional Volume Slicing + GNN TomoGraphView (ours) 8 0.7336 0.9678 0.6780 0.7878 0.9600 0.8399 0.8279
16 0.6907 0.9666 0.7244 0.7479 0.9585 0.8389 0.8212
24 0.7084 0.9655 0.7527 0.7893 0.9634 0.8441 0.8372
4.5 Impact of Volume Slice Spacing
@@
A closer examination of the results obtained with the proposed TomoGraphView framework, as shown in Table 2, reveals distinct dataset-specific trends. For head-neck, kidney, liver, and soft-tissue tumors, the best performance is achieved when using the highest number of views. In contrast, for breast and brain tumor datasets, configurations with $8$ views yield superior results compared to those with a greater number of views. To better understand these patterns, we hypothesize that the observed performance differences with respect to the number of views are influenced by dataset-specific imaging characteristics, particularly the voxel spacing along the axial dimension. As illustrated in Figure 5 (top), datasets with larger axial voxel spacing, such as soft-tissue, head-neck, and kidney, appear to benefit more from an increased number of views. This effect is reflected in the AUROC improvements observed when increasing the number of views from $8$ to $24$ in soft-tissue tumors (0.8399 vs. 0.8441), head and neck tumors (0.6780 vs. 0.7527), and kidney tumors (0.7878 vs. 0.7893).
@@
Figure 5: Box plot visualization of z-spacings (axial direction) across our utilized oncology datasets (top), and impact of volume slice spacings on the optimal number of views in terms of Area Under the Receiver Operating Characteristic Curve (AUROC) performance (bottom).
@@
We found two effects: First, the overall performance decreased from the isotropic to the anisotropic setting, which is an expected consequence of information loss and resampling artifacts. Second, and more importantly, for the anisotropic setting, the performance across different view counts now follows the same trend as in high z-spacing datasets, such as soft-tissue, head-neck, or kidney, where increasing the number of views consistently improves results, as illustrated in Table 2. This finding suggests that additional views are more suitable for compensating for the loss of information introduced by anisotropic resampling.
4.6 Graph Topology Ablation
@@
In this experiment, we conduct a graph topology ablation study, comparing different graph topologies to assess the impact of graph connectivity on our proposed TomoGraphView framework. In total, we compare five distinct graph topologies, which we refer to as spherical, uniform, linear decay, inverse, and inverse-square. More specifically, uniform, linear decay, inverse, and inverse-square topologies share the same graph connectivity (i.e., complete graph), but show differences in the edge-weights for node cross-connections (i.e., connections to non-immediate neighbors after triangulation as indicated in Section 3.3). The corresponding edge weights are determined according to the node distance in hops as indicated by Equation 5 and depicted in Figure 6 (top). In contrast, the spherical scheme does not use any node cross-connections, thus only connects nodes adjacent in triangulation (see Section 3.3), effectively representing a localized graph structure. Hence, the spherical topology results in a substantially sparser graph representation compared to the fully connected variants.
@@
We conducted the graph topology experiments across all datasets and report the average AUROC performance using configurations with 24 nodes, as previous experiments in Section 4.4 demonstrated this to be the overall best-performing setting. Additionally, a larger number of nodes enables a more informative comparison, since the addition of edges has a greater relative effect in denser graphs. The corresponding results are presented in Figure 6 (bottom). Several key observations emerge from these experiments. First, the spherical topology achieves comparable performance to the fully connected uniform topology, despite relying on substantially fewer edges. This highlights the efficiency and relevance of the local spherical connectivity pattern. Second, introducing cross-connections between all nodes improves performance only when employing smooth distance-based weighting schemes such as inverse and inverse-square, which effectively expand the receptive field by aggregating information from diverse perspectives. Among these, the inverse weighting outperforms the inverse-square variant. In contrast, the linear-decay weighting scheme, although also expanding the receptive field, leads to a decline in AUROC performance. This finding highlights the importance of using decaying weighting functions that place a stronger relative emphasis on local connections (i.e., one-hop neighbors) compared to global ones (i.e., cross-connections). Detailed results, presented on a per-dataset basis for each of the different graph topologies, can be found in the appendix (D).
@@
Figure 6: Graph Topology Analysis: Edge weighting schemes for uniform, inverse, inverse-square, and linear-decay according to the node hop distance (top). Graph topology results in terms of Area Under the Receiver Operating Characteristic Curve (AUROC) performance for each of the edge weighting schemes (bottom).
4.7 Slice-wise Feature Aggregation Methods
@@
In this section, we compare TomoGraphView with two established slice-wise aggregation methods from the literature: the long short-term memory (LSTM) model, which interprets a 3D volume as a sequence of consecutive slice embeddings, and the medical slice transformer (MST), which employs an attention mechanism to fuse intra-slice information. To ensure a fair comparison, the number of trainable parameters is fixed to 100100k across all models. Additionally, we evaluate both LSTM- and MST-based aggregation when combined with our proposed omnidirectional slicing strategy, thereby disentangling the effects of the aggregation mechanism from those of slice selection.
@@
We start by evaluating the LSTM and MST approaches in their original formulation using consecutive axial slices (2D-axial+). The results show that the best-performing MST outperforms the best-performing LSTM for both evaluation metrics, with AUROC values of $0.7768$ vs. $0.7446$ and MCC values of $0.3995$ vs. $0.3365$, highlighting the strong aggregation capability of the MST. Furthermore, this performance advantage is consistent across all view configurations, with the MST surpassing the LSTM in both evaluation metrics. Nevertheless, both methods fall short of the proposed TomoGraphView, which achieves an AUROC of $0.8372$ and an MCC of $0.5191$.
@@
Admittedly, comparing LSTM and MST in their original 2D-axial+ slicing configuration to TomoGraphView is only partially fair, as differences arise not only from the aggregation mechanism but also from the contextual information available in the input slices. To address this, we further extend the evaluation by applying omnidirectional slicing to both LSTM and MST. Two key observations emerge: (1) MST continues to outperform LSTM across all view configurations, and (2) both models experience substantial performance improvements. Specifically, for LSTM, omnidirectional slicing increases AUROC from $0.7446$ to $0.7863$ and MCC from $0.3365$ to $0.4444$. Similarly, for MST, AUROC improves from $0.7768$ to $0.8198$ and MCC from $0.3967$ to $0.4732$. These results highlight the clear benefit and broad applicability of the proposed omnidirectional volume slicing strategy. While both baselines show notable gains, TomoGraphView still achieves the highest overall performance.
@@
A more detailed, dataset-wise comparison is presented in Figure 7, which reports AUROC (left) and MCC (right) scores for $24$ views. Comprehensive numeric results across datasets, models, and view configurations are provided in Appendix Table 5. Several trends can be observed: in terms of AUROC, TomoGraphView performs comparably to baseline methods on the breast, brain, and soft-tissue tumor datasets, but clearly surpasses them on the head-neck, kidney, and liver tumor datasets. A similar pattern holds for MCC, where TomoGraphView consistently outperforms all competing methods, except for the kidney tumor dataset, where both MST variants slightly exceed its performance. Overall, these results underscore the strong capability of TomoGraphView for slice-wise feature aggregation and establish it as a competitive alternative to existing approaches.
Figure 7: Detailed results comparing slice-wise aggregation methods, including our TomoGraphView framework, across datasets for Area Under the Receiver Operating Characteristic Curve (AUROC) and Matthews correlation coefficient (MCC) performance, indicated as the average for 5-folds.
@@
Table 3: Quantitative results comparing slice-wise feature aggregation methods across their original slicing strategy (2D-axial+) and our proposed omnidirectional volume slicing, against our TomoGraphView framework. Here, LSTM corresponds to Long Short-Term Memory and MST to Medical Slice Transformer. Methods are compared across different numbers of views and performance metrics, such as the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Matthews correlation coefficient (MCC). Results indicate the average across all six datasets and 5-folds.
@@
Model Volume Slicing Views AUROC MCC
LSTM 2D-axial+ 8 $0.7305$ $0.3101$
16 $0.7330$ $0.3129$
24 $0.7446$ $0.3365$
LSTM Omnidirectional 8 $0.7764$ $0.4225$
16 $0.7744$ $0.4178$
24 $0.7863$ $0.4444$
MST 2D-axial+ 8 $0.7583$ $0.3442$
16 $0.7768$ $0.3967$
24 $0.7746$ $0.3995$
MST Omnidirectional 8 $0.8083$ $0.4691$
16 $0.8171$ $0.4732$
24 $0.8198$ $0.4530$
TomoGraphView Omnidirectional 8 0.8279 $0.4891$
16 $0.8212$ 0.5096
24 0.8372 0.5191
4.8 Comparative Analysis - 33D Pretrained Models

### v2

@@ 4.3 Impact of Volume Slicing Strategy
Incorporating additional complementary perspectives, as in the 2.5D slicing strategy, leads to a clear improvement in performance. For example, the average AUROC computed across all datasets increases from 0.7335 for the 2D-axial+ configuration (three slices) to 0.7681 for the 2.5D setup (three slices). This trend extends to the 2.5D+ variant, which consistently outperforms its 2D-axial+ counterpart for the same number of views. While the performance differences are more pronounced between $3$, 88, and $16$ views, the gap narrows at $24$ views, suggesting that incorporating diverse perspectives is particularly beneficial when only a limited number of slices are available.
@@
Beyond canonical plane slicing, our proposed omnidirectional slicing strategy, which integrates both canonical and non-canonical orientations, further improves downstream performance. On average, across all evaluated configurations, the omnidirectional approach consistently outperforms both the 2D-axial+ and 2.5D+ methods. Furthermore, for our volumetric slicing strategy, we observe a general trend across datasets. Increasing the number of views typically leads to higher performance. This relationship remains consistent for the omnidirectional approach, with the sole exception of the kidney dataset, highlighting a clear positive correlation between the number of views and overall model performance. Furthermore, although the proposed omnidirectional slicing strategy does not achieve the highest performance on every individual dataset, it consistently surpasses its 2D-axial+ and 2.5D+ counterparts for the same number of views when considering the mean performance across all datasets. This demonstrates the method’s robust and, in most cases, superior capability in capturing versatile volumetric information relevant for downstream classification.
4.4 Impact of Graph Neural Network Prediction Head
@@
Following the benchmarking results in Section 4.3, where our omnidirectional slicing strategy showed the strongest performance, we now turn to aggregating slice-level representations using the spherical mesh–based graph topology and its associated graph-based feature aggregation strategy introduced in Section 3.3. To evaluate the impact of the GNN-based prediction head, we compare two models: MLP – Omnidirectional Slicing and GNN – Omnidirectional Slicing, the latter constituting our full TomoGraphView framework. Both methods operate on identical omnidirectional slice inputs, ensuring that any performance differences arise solely from the aggregation method rather than the underlying representations. The corresponding results are summarized in Table 1.
@@
Our findings support the advantage of propagating information across slices via a meaningful underlying graph structure. Specifically, when comparing the average performance of MLP - Omnidirectional Slicing against our TomoGraphView framework (GNN - Omnidirectional Slicing) across all six datasets, our proposed framework consistently outperforms the MLP for every view configuration, ultimately increasing from $0.8154$ to $0.8372$ in AUROC performance for the best performing model in each case. Notably, our TomoGraphView framework achieves superior performance across nearly all dataset–view combinations, with the exception of Head-Neck at $8$ views and Soft-Tissue at $24$ views. These findings strongly highlight the effectiveness of the graph neural network-based feature aggregation approach over the MLP baseline method.
In summary, although our TomoGraphView framework consistently improves the overall average performance across all experiments presented in Table 1, it does not always achieve the single best score on every dataset, though it generally remains very close. We attribute these variations to reflect underlying dataset-specific characteristics, which we discuss further in Section 5.2.
4.5 TomoGraphView - Impact of Volume Slice Spacing
@@
A closer examination of the results obtained with the proposed TomoGraphView framework, as shown in Table 1, reveals distinct dataset-specific trends. For head-neck, kidney, liver, and soft-tissue tumors, the best performance is achieved when using the highest number of views. In contrast, for breast and brain tumor datasets, configurations with $8$ views yield superior results compared to those with a greater number of views. To better understand these patterns, we hypothesize that the observed performance differences with respect to the number of views are influenced by dataset-specific imaging characteristics, particularly the voxel spacing along the axial dimension. As illustrated in Figure 5 (left), datasets with larger axial voxel spacing, such as soft-tissue, head-neck, and kidney, appear to benefit more from an increased number of views. This effect is reflected in the AUROC improvements observed when increasing the number of views from $8$ to $24$ in soft-tissue tumors (0.8399 vs. 0.8441), head and neck tumors (0.6780 vs. 0.7527), and kidney tumors (0.7878 vs. 0.7893).
@@
Figure 5: Box plots illustrating the distribution of axial z-spacings across the oncology datasets used in this study (left), and the effect of slice spacing on the optimal number of views for TomoGraphView, shown for the brain tumors dataset using Area Under the Receiver Operating Characteristic Curve (AUROC) as the performance metric (right).
@@
As visualized in Figure 5 (right), we observe two effects: First, the overall performance decreased from the isotropic to the anisotropic setting, which is an expected consequence of information loss and resampling artifacts. Second, and more importantly, for the anisotropic setting, the performance across different view counts now follows the same trend as in high z-spacing datasets, such as soft-tissue, head-neck, or kidney, where increasing the number of views consistently improves results, as illustrated in Table 1. This finding suggests that additional views are more suitable for compensating for the loss of information introduced by anisotropic resampling.
4.6 TomoGraphView - Impact of Graph Topology
@@
We continue by studying the impact of the underlying graph topology, comparing different graph structures to assess the effect of graph connectivity on our proposed TomoGraphView framework. In total, we compare five distinct graph topologies, which we refer to as spherical, uniform, linear decay, inverse, and inverse-square. More specifically, uniform, linear decay, inverse, and inverse-square topologies share the same graph connectivity (i.e., complete graph), but show differences in the edge-weights for node cross-connections (i.e., connections to non-immediate neighbors after triangulation as indicated in Section 3.3). The corresponding edge weights are determined according to the node distance in hops as indicated by Equation 5 and depicted in Figure 6 (top). In contrast, the spherical scheme does not use any node cross-connections, thus only connects nodes that are adjacent in the Delaunay triangulation (see Section 3.3), effectively representing a localized graph structure. Hence, the spherical topology results in a substantially sparser graph representation compared to the fully connected variants.
We conducted the graph topology experiments across all datasets and report the average AUROC performance using configurations with 24 nodes, as previous experiments in Section 4.4 demonstrated this to be the overall best-performing setting for TomoGraphView. Additionally, a larger number of nodes enables a more informative comparison, since the addition of edges has a greater relative effect in denser graphs. The corresponding results are presented in Figure 6 (bottom).
@@
Several key observations emerge from these experiments. First, the spherical topology achieves comparable performance to the fully connected uniform topology, despite relying on substantially fewer edge connections. This highlights the efficiency and relevance of the local spherical graph connectivity pattern. Second, introducing cross-connections between all nodes improves performance only when employing smooth distance-based weighting schemes such as inverse and inverse-square, which effectively expand the receptive field by aggregating information from diverse perspectives. Among these, the inverse weighting outperforms the inverse-square variant. In contrast, the linear-decay weighting scheme, although also expanding the receptive field, leads to a decline in AUROC performance. This finding highlights the importance of using decaying weighting functions that place a stronger relative emphasis on local connections (i.e., one-hop neighbors) compared to global ones (i.e., node cross-connections). Detailed results, presented on a per-dataset basis for each of the different graph topologies, are provided in the D.
@@
Figure 6: Edge-weighting strategies based on node hop distance, including uniform, inverse, inverse-square, and linear-decay schemes (top). Corresponding graph topology performance for 24 views is shown using Area Under the Receiver Operating Characteristic Curve (AUROC) as the evaluation metric for each weighting scheme (bottom).
4.7 Slice-wise Feature Aggregation Method Benchmark
@@
In this section, we compare TomoGraphView with two established slice-wise aggregation methods: a long short-term memory (LSTM) approach, which interprets a 33D volume as a sequence of consecutive slice embeddings, and a transformer-based approach, represented by the medical slice transformer (MST) [muller2025medical], which employs an attention mechanism to fuse intra-slice information. To ensure a fair comparison with previous experiments, the number of trainable parameters is again fixed to 100100k across all models. Additionally, we evaluate both LSTM- and transformer-based (MST) aggregation when combined with our proposed omnidirectional slicing strategy, thereby disentangling the effects of the aggregation mechanism from those of slice selection.
@@
We start by evaluating the LSTM and MST approaches in their original formulation using consecutive axial slices (i.e., 2D-axial+). The results show that the best-performing view setting ($16$) in the MST approach outperforms the best-performing view setting ($24$) in the LSTM approach, for both evaluation metrics, with AUROC values of $0.7768$ vs. $0.7446$ and MCC values of $0.3995$ vs. $0.3365$, highlighting the strong aggregation capability of the MST. Furthermore, this performance advantage is even observable across all view configurations, with the MST again surpassing the LSTM in both evaluation metrics. Nevertheless, both methods fall short of our proposed TomoGraphView framework, which overall achieves an AUROC of $0.8372$ and an MCC of $0.5191$.
@@
Admittedly, comparing LSTM and MST in their original 2D-axial+ slicing configuration to TomoGraphView is only partially fair, as differences may not only arise from the aggregation mechanism but also from the contextual information available in the input slices. To address this, we further extend the evaluation by applying omnidirectional slicing to both LSTM and MST. Two key observations emerge: (I) MST continues to outperform LSTM across all view configurations, and (II) both models experience substantial performance improvements. Specifically, for LSTM, omnidirectional slicing increases AUROC from $0.7446$ to $0.7863$ and MCC from $0.3365$ to $0.4444$. Similarly, for MST, AUROC improves from $0.7768$ to $0.8198$ and MCC from $0.3967$ to $0.4732$. These results highlight the clear benefit and broad applicability of our proposed omnidirectional volume slicing strategy. Remarkably, while both baselines show notable gains, TomoGraphView still achieves the highest overall performance.
@@
A more detailed, dataset-wise comparison is presented in Figure 7, which reports AUROC (left) and MCC (right) scores for $24$ views. Comprehensive numeric results across datasets, models, and view configurations are provided in Appendix Table 4. Several trends can be observed: in terms of AUROC, TomoGraphView matches the performance of the best baseline method on the breast, brain, and soft-tissue tumor datasets, but clearly surpasses them on the head-neck, kidney, and liver tumor datasets. A similar pattern holds for MCC, where TomoGraphView consistently outperforms all competing methods, except for the kidney tumor dataset. Here, both MST variants show superior performance. Nevertheless, the overall results underscore the strong capability of TomoGraphView for slice-wise feature aggregation, establishing it as both a competitive and compelling alternative to existing slice-wise approaches.
@@
Figure 7: Detailed results comparing slice-wise aggregation methods: long short-term memory (LSTM), medical slice transformer (MST), and TomoGraphView. Besides their original slicing strategy (2D-axial+), LSTM, and MST are also evaluated using omnidirectional volume slicing. All methods are evaluated across all six datasets using the Area Under the Receiver Operating Characteristic Curve (AUROC) on the left and Matthews correlation coefficient (MCC) on the right as performance metrics. Bars indicate the average performance across five folds.
Table 2: Quantitative results comparing a long short-term memory (LSTM) approach, medical slice transformer (MST), and our TomoGraphView framework for slice-wise feature aggregation performance. Besides their original slicing strategy (2D-axial+), LSTM and MST are also evaluated using omnidirectional volume slicing. Methods are compared across different numbers of views and performance metrics, including the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Matthews correlation coefficient (MCC). Results are presented as the average across all six datasets and five folds.
@@
Model Volume Slicing Views AUROC MCC LSTM 2D-axial+ 8 $0.7305$ $0.3101$ 16 $0.7330$ $0.3129$ 24 $0.7446$ $0.3365$ LSTM Omnidirectional 8 $0.7764$ $0.4225$ 16 $0.7744$ $0.4178$ 24 $0.7863$ $0.4444$ MST 2D-axial+ 8 $0.7583$ $0.3442$ 16 $0.7768$ $0.3967$ 24 $0.7746$ $0.3995$ MST Omnidirectional 8 $0.8083$ $0.4691$ 16 $0.8171$ $0.4732$ 24 $0.8198$ $0.4530$ TomoGraphView Omnidirectional 8 0.8279 $0.4891$ 16 $0.8212$ 0.5096 24 0.8372 0.5191
4.8 Comparative Analysis with 33D Pretrained Models